# Disentangling modified gravity from a dark force with gravitational redshift

### v1

@@ IV.1 The general approach
In order to quantify the degeneracy between the GBD and CQ scenarios and assess the ability of gravitational redshift to break it, we use mgcamb Zhao et al. (2009); Hojjati et al. (2011); Zucca et al. (2019); Wang et al. (2023), a modified gravity patch of the Boltzmann code camb Lewis et al. (2000); Howlett et al. (2012). We employ Cobaya Torrado and Lewis (2019, 2021) to carry out an MCMC analysis of the mock data set consisting of the observables described in Sec. III. As a first step, we generate the mock data for a given fiducial model involving one kind of modification only, for example CQ with $\beta_{*2}=1$ and fixed mass $m_{*2}$. We then perform a fit of the data allowing for both $\beta_{*1}$ and $\beta_{*2}$ to vary, fixing $m_{*1}=m_{*2}$. [^4] Note that by simultaneously varying both $\beta_{*1}$ and $\beta_{*2}$, we are not claiming that we live in a universe where both gravity modifications and a dark fifth force are present. This approach is necessary to determine which models can be constrained by the data, without arbitrarily excluding one of the modifications. The resulting constraints will either indicate the presence of a degeneracy or the lack thereof: if the two configurations $\{\beta_{*1}=0,\beta_{*2}\neq 0\}$ and $\{\beta_{*1}\neq 0,\beta_{*2}=0\}$ provide an equally good fit, we can conclude that the data cannot distinguish between the two scenarios. On the other hand, if the case $\{\beta_{*1}\neq 0,\beta_{*2}=0\}$ is excluded, the two kinds of modifications can be disentangled from one another, breaking the degeneracy.
For each data vector, we perform two types of analysis to assess the degeneracy between the two models: one where we do not include the dipole in the clustering data in Eq. (15), i.e. there are no constraints from gravitational redshift, and another one where the dipole is included. We build a multivariate Gaussian likelihood for the combination of the three observables and adopt wide uniform priors on the free parameters.
@@ IV.2.1 Galaxy clustering
This includes the usual even multipoles of RSD and importantly, also the dipole $\xi^{\mathrm{BF}}_{1}$. Note that we only consider the hexadecapole in the total population of galaxies (labelled T), as Eq. (12) does not contain a dependence on population-specific biases so that considering different galaxy samples does not provide additional information. We compute the multipoles in Fourier space from power spectra generated using mgcamb, and perform the transform to configuration space using the FFTLog method Hamilton (2000).
Since the dipole is too small to be measured from current data Gaztanaga et al. (2017), we perform the forecasts assuming a SKA2-like survey, for which the dipole is expected to reach a signal-to-noise ratio of 80 Castello et al. (2023). We consider the specifications of Ref. Bull (2016) for the number density, sky coverage and galaxy bias. The separations considered in this work range from $d=20\,h^{-1}$pc to $d=160\,h^{-1}$pc, an interval where nonlinear corrections were found to be negligible for the dipole Bonvin et al. (2023). Furthermore, we assume a bias difference of 1 between the bright and faint galaxy populations, in agreement with that measured in Ref. Gaztanaga et al. (2017). More details on the modelling of the galaxy bias and the magnification bias can be found in Appendix B.
@@ IV.2.2 Gravitational lensing
For gravitational lensing, we consider the specifications of the Dark Energy Survey (DES) Year 1 Abbott et al. (2018); Krause et al. (2017) and compute the shear-shear correlations and the galaxy-galaxy lensing correlations using mgcamb. Since mgcamb is based on linear perturbation theory, we impose the “aggressive” scale cut implemented in Ref. Zucca et al. (2019) to effectively remove the data in the nonlinear regime. This is more restrictive than the cut used in the DES analysis for the $\Lambda$CDM model Abbott et al. (2018), since in our case it would be incorrect to follow the same approach and assume the validity of GR to model the observables on mildly nonlinear scales. All nuisance parameters, including intrinsic alignment, lens photo-z shift, source photo-z shift and shear calibration are fixed according to the DES Year 1 standard values Abbott et al. (2018); Krause et al. (2017).
@@ V.1 Breaking the degeneracy
In Fig. 1, we show the marginalized constraints on $\beta_{*1}$ and $\beta_{*2}$ for a CQ fiducial model with $\beta_{*2}=1$, $\lambda_{*}=10$ Mpc, a combination that is not excluded by current data, see Ref. Wang et al. (2023). The blue contours correspond to the analysis without the dipole, i.e. without gravitational redshift, while the red contours include it. As expected, when the dipole is not present, we obtain a perfect elliptic degeneracy between $\beta_{*1}$ and $\beta_{*2}$, which matches the analytical expectation discussed in Sec. II.2. This means that even though no gravity modifications were included in the mock data vector, the data (RSD, weak lensing, and CMB) are equally well described using either the (correct) CQ model or the GBD one.
In the GBD case, the best fit is around $\beta_{*1}=0.84$ and not $\beta_{*1}=1$. This reflects the fact that the impact of a dark fifth force with coupling strength $\beta_{*2}=1$ on the growth of cosmic structure can be mimicked by a modification of gravity with smaller coupling $\beta_{1}=(\rho_{c}/\rho)(\delta_{c}/\delta)^{1/2}\beta_{2}\simeq(\rho_{c}/\rho)^{3/2}\beta_{2}\simeq 0.84\,\beta_{2}$, as can be seen from Eqs. (3)-(4) with our choice of fiducial cosmological parameters. This degeneracy has important consequences when jointly analyzing RSD, weak lensing and CMB data: if CDM is subject to additional interactions, such modifications in the dark sector could be incorrectly interpreted as evidence for modified gravity, even though GR remains valid.
The inclusion of gravitational redshift into the analysis decisively breaks the degeneracy between the two models. Indeed, we can clearly see that the blue contours exclude the case $\beta_{*2}=0$, indicating that a pure modification of gravity no longer fits the data. Even with the dipole, a large portion of the parameter space is still allowed, implying that models with both a dark fifth force and a modification of gravity are not excluded. However, such scenarios involving both kinds of modifications would be disfavoured according to the Occam’s razor.
@@ V.2 Varying the coupling strength
Figure 2: The $1\sigma$ confidence regions for four different CQ couplings $\beta_{*2}$. For all fiducial models (indicated by stars) we set $\lambda_{*}=10$ Mpc. All other parameters, except for $\beta_{*1}$ and $\beta_{*2}$, have been fixed to their fiducial values.
@@ V.4 The range of validity of our results
All the results presented so far concern GR models with a dark fifth force acting on CDM. Since the situation is symmetric, the same conclusions also apply in the case of GBD modified gravity with no CDM interactions. In this case, the elliptic degeneracy without gravitational redshift will be such that mock data generated by a GBD model with $\beta_{*1}=1$ can be well described by a CQ scenario with $\beta_{2}\simeq(\rho_{c}/\rho)^{-3/2}\simeq 1.19$. Including gravitational redshift allows us to distinguish between the two scenarios down to $\beta_{*1}\simeq 0.7$ for $\lambda_{*}=10$ Mpc.
@@ VI Conclusions
Combining RSD with gravitational lensing is generally regarded as the optimal way to test for deviations from GR. Since RSD probe the evolution of structure, while gravitational lensing is sensitive to the geometry of the Universe, their combination would test for the gravitational slip Amendola et al. (2008); Daniel et al. (2009); Zhao et al. (2010); Song et al. (2011); Motta et al. (2013), which in a modified gravity scenario is due to an additional degree of freedom mediating a force on all non-relativistic matter. However, as we have shown in this paper, these two observables cannot distinguish a modification of gravity from additional forces acting on CDM. By performing an MCMC analysis, we have demonstrated that mock data generated by a coupled quintessence model with a fifth force acting on CDM is equally well fitted by a GBD modified gravity scenario. This means that, in the future, we may wrongly claim to have discovered a deviation from GR on cosmological scales, whereas in reality the modifications are due to new interactions in the dark sector (or viceversa).
@@ Acknowledgements.
Appendix A Generalized Brans Dicke and Coupled Quintessence scenarios
@@
We adopt the line element in Eq. (1) and work under the quasi-static approximation, neglecting time derivatives of the metric and the field perturbations over spatial ones Mirpoorian et al. (2023). For GBD, this leads to the following sets of equations in Fourier space on sub-horizon scales:
@@
Here, $\Box\equiv\nabla^{\mu}\nabla^{\mu}$ and the scalar field coupling strength is given by $\beta_{1}=A_{,\phi}/{A}$, where a comma indicates a derivative, and $\tilde{\beta}_{1}=\beta_{1}/\sqrt{8\pi G}$.
@@ Appendix B Galaxy clustering survey specifications
where P indicates the bright (B) or faint (F) galaxy population. We let the parameters $b_{\mathrm{P,1}}$ and $b_{\mathrm{P,2}}$ free with fiducial values $0.554$ and $0.873$ respectively, and set $\Delta b=1$, in agreement with the bias difference measured in Ref. Gaztanaga et al. (2017).

### v2

@@ IV.1 The general approach
In order to quantify the degeneracy between the GBD and CQ scenarios and assess the ability of gravitational redshift to break it, we use mgcamb Zhao et al. (2009); Hojjati et al. (2011); Zucca et al. (2019); Wang et al. (2023), a modified gravity patch of the Boltzmann code camb Lewis et al. (2000); Howlett et al. (2012). We employ Cobaya Torrado and Lewis (2019, 2021) to carry out an MCMC analysis of the mock dataset consisting of the observables described in Sec. III.
We generate mock data for a given fiducial model where only one kind of modification is present, for example CQ with $\beta_{*2}=1$ and a given mass $m_{*2}$. However, when we perform a fit of the data, we treat both $\beta_{*1}$ and $\beta_{*2}$ as free parameters, fixing $m_{*1}=m_{*2}$. This does not mean that both modifications are likely to occur at the same time in our Universe, but it is the correct procedure to properly quantify the degeneracy between the two models and determine whether the input fiducial model can be recovered. The resulting constraints will either indicate the presence of a degeneracy or the lack thereof: if the two configurations $\{\beta_{*1}=0,\beta_{*2}\neq 0\}$ and $\{\beta_{*1}\neq 0,\beta_{*2}=0\}$ provide an equally good fit, we can conclude that the data cannot distinguish between the two scenarios. On the other hand, if the case $\{\beta_{*1}\neq 0,\beta_{*2}=0\}$ is excluded, the two kinds of modifications can be disentangled from one another, breaking the degeneracy.
For each mock data vector built according to the specifications in Sec. IV.2, we perform two types of analysis to assess the degeneracy between the two models: one where we do not include the dipole in the clustering data in Eq. (16), i.e. there are no constraints from gravitational redshift, and another one where the dipole is included. We build a multivariate Gaussian likelihood for the combination of the three observables and adopt wide uniform priors on the free parameters.
@@ IV.2.1 Galaxy clustering
This includes the usual even multipoles of RSD and importantly, also the dipole $\xi^{\mathrm{BF}}_{1}$. Note that we only consider the hexadecapole in the total population of galaxies (labeled T), as Eq. (13) does not contain a dependence on population-specific biases and thus considering different galaxy samples does not provide additional information. We compute the multipoles in Fourier space from power spectra generated using mgcamb and perform the transform to configuration space using the FFTLog method (Fast Fourier Transform with Logarithmically spaced periodic sequence) Hamilton (2000).
Since the dipole is too small to be measured from current data Gaztanaga et al. (2017), we perform the forecasts assuming a SKA2-like survey, for which the dipole is expected to reach a signal-to-noise ratio of 80 Castello et al. (2024). We consider the specifications of Ref. Bull (2016) for the number density, sky coverage, and galaxy bias. The separations considered in this work range from $d=20\,h^{-1}$pc to $d=160\,h^{-1}$pc, an interval where nonlinear corrections were found to be negligible for the dipole Bonvin et al. (2023). Furthermore, we assume a bias difference of 1 between the bright and faint galaxy populations, in agreement with that measured in Ref. Gaztanaga et al. (2017). More details on the modelling of the galaxy bias and the magnification bias can be found in Appendix B.
@@ IV.2.2 Gravitational lensing
For gravitational lensing, we consider the specifications of the Dark Energy Survey (DES) Year 1 Abbott et al. (2018); Krause et al. (2017) and compute the shear-shear correlations and the galaxy-galaxy lensing correlations using mgcamb. Since mgcamb is based on linear perturbation theory, we impose the “aggressive” scale cut implemented in Ref. Zucca et al. (2019) to effectively remove the data in the nonlinear regime. This is more restrictive than the cut used in the DES analysis for the $\Lambda$CDM model Abbott et al. (2018), since in our case it would be incorrect to follow the same approach and assume the validity of GR to model the observables on mildly nonlinear scales. All nuisance parameters, including intrinsic alignment, lens photo-z shift, source photo-z shift, and shear calibration are fixed according to the DES Year 1 standard values Abbott et al. (2018); Krause et al. (2017).
@@ V.1 Breaking the degeneracy
In Fig. 1, we show the marginalized constraints on $\beta_{*1}$ and $\beta_{*2}$ for a CQ fiducial model with $\beta_{*2}=1$, $\lambda_{*}=10$ Mpc, a combination that is not excluded by current data, see Ref. Wang et al. (2023). The blue contours correspond to the analysis without the dipole, i.e. without gravitational redshift, while the red contours include it. As expected, when the dipole is not present, we obtain a perfect elliptic degeneracy between $\beta_{*1}$ and $\beta_{*2}$, which is consistent with the analytical expectation in Eq. (5) for our chosen time evolution in Eq. (7). This means that even though no gravity modifications were included in the mock data vector, the data (RSD, weak lensing, and CMB) are equally well described using either the (correct) CQ model or the GBD one.
In the GBD case, the best fit is around $\beta_{*1}=0.84$ and not $\beta_{*1}=1$. This reflects the fact that the impact of a dark fifth force with coupling strength $\beta_{*2}=1$ on the growth of cosmic structure can be mimicked by a modification of gravity with smaller coupling $\beta_{1}=(\rho_{c}/\rho)(\delta_{c}/\delta)^{1/2}\beta_{2}\simeq(\rho_{c}/\rho)^{3/2}\beta_{2}\simeq 0.84\,\beta_{2}$, as can be seen from Eqs. (3)-(4) with our choice of fiducial cosmological parameters. This degeneracy has important consequences when jointly analyzing RSD, weak lensing, and CMB data: if CDM is subject to additional interactions, such modifications in the dark sector could be incorrectly interpreted as evidence for modified gravity, even though GR remains valid.
The inclusion of gravitational redshift into the analysis decisively breaks the degeneracy between the two models. Indeed, we can clearly see that the blue contours exclude the case $\beta_{*2}=0$, indicating that a pure modification of gravity no longer fits the data. Even with the dipole, a large portion of the parameter space is still allowed, implying that models with both a dark fifth force and a modification of gravity are not excluded. However, such scenarios involving both kinds of modifications would be disfavoured according to the Occam’s razor. An alternative representation of the constraints in polar coordinates is presented in Appendix C.
@@ V.2 Varying the coupling strength
It should be noted that the breaking of the degeneracy between $\beta_{*1}$ and $\beta_{*2}$ does not arise from the total dipole in Eq. (III.1), but only from the part induced by the gravitational redshift, namely, the $P_{\delta\Psi}$ term. It is only this specific term that carries the additional information to distinguish between the two models and determines the lower bound in $\beta_{*2}$ for the dipole to be effective.
Figure 2: The $1\sigma$ confidence regions for four different fiducial values of the CQ coupling $\beta_{*2}$. For all fiducial models (indicated by stars) we set $\lambda_{*}=10$ Mpc. All parameters except for $\beta_{*1}$ and $\beta_{*2}$ have been fixed to their fiducial values.
@@ V.4 The range of validity of our results
All the results presented so far concern GR models with a dark fifth force acting on CDM. Since the situation is symmetric, the same conclusions also apply in the case of GBD modified gravity with no CDM interactions. In this case, the elliptic degeneracy without gravitational redshift will be such that mock data generated by a GBD model with $\beta_{*1}=1$ can be well described by a CQ scenario with $\beta_{*2}\simeq(\rho_{c}/\rho)^{-3/2}\simeq 1.19$. Including gravitational redshift allows us to distinguish between the two scenarios down to $\beta_{*1}\simeq 0.7$ for $\lambda_{*}=10$ Mpc.
@@ VI Conclusions
Combining RSD with gravitational lensing is generally regarded as the optimal way to test for deviations from GR. Since RSD probes the evolution of structure, while gravitational lensing is sensitive to the geometry of the Universe, their combination would test for the gravitational slip Amendola et al. (2008); Daniel et al. (2009); Zhao et al. (2010); Song et al. (2011); Motta et al. (2013), which in a modified gravity scenario is due to an additional degree of freedom mediating a force on all non-relativistic matter. However, as we have shown in this paper, these two observables cannot distinguish a modification of gravity from additional forces acting on CDM. By performing an MCMC analysis, we have demonstrated that mock data generated by a coupled quintessence model with a fifth force acting on CDM is equally well fitted by a GBD modified gravity scenario. This means that, in the future, we may wrongly claim to have discovered a deviation from GR on cosmological scales, whereas in reality the modifications are due to new interactions in the dark sector (or vice versa).
@@
For the purpose of this demonstration, we have considered a scenario where the time evolution of both modifications is exactly the same. This was intentionally chosen so that the two models could be treated on an equal footing and would yield the very same impact on the growth of cosmic structure. A different time evolution could in principle partly break the degeneracy obtained with RSD data, but, since this would not be known a priori when analyzing real data, we consider scenarios where the two models can exactly mimic each other.
@@ Acknowledgements.
Appendix A Generalized Brans-Dicke and coupled quintessence scenarios
@@
We adopt the line element in Eq. (1) and work under the quasi-static approximation, neglecting time derivatives of the metric and the field perturbations over spatial ones Mirpoorian et al. (2023). For GBD, this leads to the following sets of equations in Fourier space on subhorizon scales:
@@
Here, $\Box\equiv\nabla^{\mu}\nabla_{\mu}$, the scalar field coupling strength is given by $\beta_{1}=A_{,\phi}/{A}$, where a comma indicates a derivative, and we have defined $\tilde{\beta}_{1}\equiv\beta_{1}/\sqrt{8\pi G}$.
@@ Appendix B Galaxy clustering survey specifications
where P indicates the bright (B) or faint (F) galaxy population. We let the parameters $b_{\mathrm{P,1}}$ and $b_{\mathrm{P,2}}$ free with fiducial values $0.554$ and $0.873$, respectively, and we set $\Delta b=1$ in agreement with the bias difference measured in Ref. Gaztanaga et al. (2017).
@@
Appendix C Constraints in polar coordinates
An alternative visualization of the constraints can be achieved by performing a change to polar coordinates,
$R=\sqrt{\beta_{*1}^{2}+\beta_{*2}^{2}}\,,\qquad\theta=\frac{\beta_{*2}}{R}\,.$ (36)
[…]
Figure 4: The $1\sigma$ and confidence regions on the polar coordinates $R$ and $\theta$. Here, the fiducial model is CQ and is specified by $\beta_{*2}=1$ and $\lambda_{*}=10$ Mpc (with $\beta_{*1}=0$).